# Wake Vision: A Large-scale, High-Quality Dataset and Benchmark Suite for TinyML Applications

### v2

@@ 3.1 Label Generation
Table 2: Accuracy on test set trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
@@
As our distillation results indicate, further data-centric filtering techniques may be needed to fully leverage the image-level training set. We experimented with tripling the model size and doubling the training time, but the difference in accuracy remained relatively consistent.
The two training sets provide a useful foundation for future work on the importance of data quality vs. quantity. Wake Vision can enable further research on data-centric TinyML [21, 22], specifically uncovering the relationship between model capacity and the sensitivity to training data quality.
3.2 Label Correction
Label errors are a challenge in computer vision [24, 4] that limit progress and obscures true performance. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision. All labels in the Open Images dataset are originally machine-generated candidate labels. Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases. Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present, but lacks the corresponding label. Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking. Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset accurately represents the model’s performance. To do so we make use of the Scale AI platform to crowd-source manual label corrections (see Appendix). With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [23]. We perform an initial exploration of these methods in appendix N. We report our estimated label error rate after label corrections versus VWW in Table 3. The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
Table 3: Estimated label error rate of VWW and Wake Vision, based on a 500 sample subset. The Wake Vision estimated error rate is broken down by set. The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different sources of their labels. The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.
Dataset Wake Vision Set Label Error Rate
Visual Wake Words Train, Val, & Test 7.8%
Wake Vision Train (Large) 15.2%
Wake Vision Train (Quality) 6.8%
Wake Vision Val & Test 2.2%
3.3 Usability
Wake Vision is available through TensorFlow Datasets [28] and HuggingFace Datasets [17] to enable easy access and use by the community. The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark. The rehosted labels are generated according to our default dataset configuration, further described in Sec. I.
3.4 Comparison to Visual Wake Words
@@
Wake Vision serves as a drop-in improvement over VWW, thanks to its size and our extensive filtering. To compare the benefits of Wake Vision against VWW, we train two identical MobileNetv2 [26] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [20]. We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6. One model uses VWW’s training set, and the other employs Wake Vision’s Quality training set. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we cross-evaluate their performance on the corresponding test sets.
@@
Table 4 shows a 0.26% improvement on VWW’s own test set, which illustrates that Wake Vision is a direct improvement over VWW, not simply a domain shift. We additionally achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging. Table 3 also shows the performance of a the Wake Vision (Combined) model described in section 3.1. This model achieves an average 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and an average 0.83% improvement on the Wake Vision test set. The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models, but shows the value of using the training sets in unison even on the VWW test set.
Table 4: Accuracy on the Wake Vision and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
Train
VWW Wake Vision (Quality) Wake Vision (Combined)
Test VWW 88.33$\pm$0.29 88.59$\pm$0.1789.34$\pm$0.02%
Wake Vision 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
3.5 Dataset Quality Vs. Size
@@
We investigate the impact of dataset quality and size of Resnet [10] style models of varying capacity. We measure dataset quality by the approximate rate of label errors. The Wake Vision (Quality) training set has an estimated error rate of around 7%. We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability. Appendix D gives the derivation for the flip probability. We also sweep the dataset size by taking a slice of Wake Vision (Quality). Each model is trained for 50,000 steps on 224x224x3 images using AdamW [20], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004. Figure 3 illustrates the experiment’s results. We observe that smaller models (leftmost figure) are more sensitive to higher error rates than large models. The smallest model’s accuracy drop is 1.3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model. In contrast, large models benefit more from big datasets and start to over fit on the smaller slices of the training data.
@@
Figure 3: Comparing the impact of dataset error rate and size on ResNet models of varying capacity. The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training. The error % indicates the expected rate of incorrectly labeled samples.
4 Fine Grain Benchmark Suite
Test sets generated directly from images posted to the internet are not reflective of real world use cases as the distribution is biased towards images people deems as worth sharing (e.g. well lit and framed). In contrast, person detection systems are deployed, and expected to work, in more challenging settings, such as low lighting. Due to this gap, a model that achieves high test accuracy may perform poorly in real-world scenarios after the design stage has ended. To address this issue, we present a benchmark suite for person detection that tests a model’s robustness in challenging settings, enabling better analysis at the design stage. Additionally, the suite measures the model’s performance across demographics to ensure it does not exhibit clear bias. The suite consists of five fine-grained benchmark sets. Each of the five fine-grained benchmarks have been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test. The benchmark determines whether a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well-lit compared to the inverse setting. Appendix C contains a case study showing how the fine grained evaluation sets can be used to design robust models. Table 5 reports the number of samples in each of the individual benchmark sets and the F1 scores of a Wake Vision and VWW model on each set.
Perceived Gender and Age
These fine grained datasets are generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [27]. Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [12]. This benchmark aims to evaluate a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias. These labels are based on a perceived gender and age representation and are not necessarily representative.
@@ 4 Fine Grain Benchmark Suite
Gender Female 684 2,157 0.93 0.89
Male 1,310 3,918 0.91 0.88
Unknown 1,612 4,940 0.77 0.78
@@
Unknown 1,299 3,837 0.71 0.74
Distance Near 5,457 16,333 0.91 0.89
Medium 2,213 6,876 0.85 0.84
Far 398 1,140 0.59 0.67
Lighting Dark 3,255 9,420 0.85 0.81
Normal 14,315 43,010 0.82 0.82
Bright 1,012 3,332 0.82 0.82
@@
No Depiction 8,583 25,802 0.87 0.85
Distance
The distance datasets test how the distance of persons in images impact the performance of ML systems. This benchmark is critical for correctly predicting a model’s deployed performance. If a person detection system is intended to recognize subjects at great distances, the system’s performance on the far away dataset will be more informative than its performance on the top-level test set. We create three datasets based on the percentage of the image the subject bounding box covers. The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
Lighting
The lighting datasets aim to test the performance of ML systems across different lighting conditions. In scenarios like security cameras, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments. We create three fine grained datasets of this type for dark, normal, and bright lighting conditions, respectively. We quantify lighting conditions by the average pixel values of images in greyscale, which is a simple but effective method for distinguishing lighting conditions [33]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
Depictions
A particularly challenging task for a person detection model is to correctly reject depictions of people. In many use cases a person detection model can not falsely trigger on a depiction. For example, a room occupancy detector that incorrectly identifies a painting on the wall as a person. This benchmark measures the models accuracy on three related sets of non-person samples: depictions of people, depictions that are not of people, and images do not contain a depiction of any kind. Depictions of people can range from photo-realistic to crude stick figures.
4.1 Benchmark Results
Table 5 compares the results of a model trained on the Wake Vision (Quality) training set to an identical model trained on VWW. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66. Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82. The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively. This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
@@ 5 Ethical Considerations
The Wake Vision dataset and benchmark suite aim to advance TinyML research by providing large-scale datasets and enabling fine-grained analysis of TinyML systems. Wake Vision can directly improve the quality of person detection systems that benefit society, such as energy-saving applications, while protecting user privacy through on-device inference. However, person detection models can also be used for malicious purposes, such as weapons targeting or large-scale surveillance.
@@
The images in Wake Vision are sourced from Flickr through Open Images[16, 14] under the CC-BY 2.0 license. While we have made efforts to ensure the images are properly licensed, we cannot guarantee the license status of each image. It is possible that some images have been uploaded without the right to distribute under the CC-BY 2.0 license. Given sufficient resources, individuals in the dataset could potentially be identified based on biometric data. The dataset may also contain offensive content, as manually verifying millions of images is infeasible.
The benchmark suite aims to ensure the fairness and robustness of person detection models. However, the demographic benchmarks could potentially be misused to create systems that classify gender and age, breaching personal privacy or perpetuating discrimination. To mitigate this risk, we have not released training sets for the fine-grained benchmark datasets.
6 Conclusions
The lack of large, high-quality datasets has been a significant bottleneck in the field of TinyML research. To address this issue, we introduced Wake Vision, a large-scale dataset and benchmark suite for person detection that surpasses the current state-of-the-art dataset, VWW, in terms of size (~100x larger), ease of use, label accuracy, and real-world relevance. These improvements make Wake Vision a drop-in replacement for VWW. Recognizing the importance of fairness across demographics and robustness to challenging settings in person detection use cases, we also released a Wake Vision benchmark suite comprising six fine-grained benchmark sets to evaluate the fairness and robustness of person detection models. Furthermore, our investigation into the trade-off between dataset quality and quantity for TinyML systems suggests that dataset quality is likely more crucial for TinyML systems compared to conventional ML systems. The Wake Vision dataset and benchmark suite aim to facilitate the development of more accurate, fair, and robust person detection models in TinyML.
7 Acknowledgements
This work was partially supported by the Google TPU Research Cloud program. We would like to thank Andrew Howard of Google Deepmind for his input and support in this effort.
Appendix A Code Repository

### v3

@@ 4.1 Training Set Evaluation
Table 4: Accuracy on test set of a MobileNetV2-0.25 trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
@@
To evaluate the performance of the two Wake Vision training sets, we train identical models on each set. We use MobileNetV2 [33] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [26]. We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6. After training, we compare each model’s performance on the common test set.
The result of this comparison is shown in Tab. 4. These results show that the Wake Vision (Quality) training set outperforms the Wake Vision (Large) training set by 4.09% test accuracy. This indicates that label quality is more important than quantity in this setting. The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
Using the two training sets in unison to train a model where Wake Vision (Large) acts as a pre-training set, and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%. The two training sets together provide a foundation for further research on data-centric TinyML [27, 28].
4.2 Wake Vision & VWW Cross Evaluation
A key challenge in advancing TinyML research is ensuring that improvements can be readily adopted by the community without requiring significant modifications to existing workflows or architectures. While novel datasets can offer better quality or scale, their practical impact is limited if they require substantial changes to model architectures, training pipelines, or deployment processes. Therefore, we designed Wake Vision to serve as a drop-in replacement for VWW.
To validate this drop-in capability and demonstrate Wake Vision’s effectiveness, we do a cross-evaluation study. We train two identical MobileNetv2 [33] models according to the same training recipe as in Sec. 4.1. One model uses VWW’s training set, and the other uses Wake Vision’s Quality training set. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we evaluate their performance on the corresponding test sets. Creftab:wv-vww-cross-val shows the result of the cross-evaluation. The Wake Vision trained model outperforms the VWW trained model with a 0.26% improvement on VWW’s own test set, which indicates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
@@ 4.2 Wake Vision & VWW Cross Evaluation
Also, we achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging. Table 3 also shows the performance of the Wake Vision (Combined) model described in Sec. 3.1. This model achieves 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and 0.83% improvement on the Wake Vision test set. The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models but shows the value of using the training sets in unison even on the VWW test set.
4.3 Comprehensive Model Architecture Evaluation
To show that the benefits of using Wake Vision over VWW are not limited to MobileNetv2 architectures, we also train several other models on both training sets, Wake Vision (Quality) and VWW, and compare their performance. For this experiment, we use a collection of models commonly used in the TinyML community, such as models from the MCUNet [24] and Micronets [2] families. Furthermore, we include models from ColabNAS [11] to provide evaluation results on even smaller models.
Figure 3: Cross-evaluation results on the Wake Vision test set.
Each model is trained on each dataset using the Adam optimizer with a learning rate of 0.001 for 243,800 steps at a batch size of 512, equivalent to 100 epochs on the Wake Vision (Quality) training set. In addition, we apply random horizontal flips rotations and use 8-bit quantization-aware training. We use the checkpoint with the best validation score for testing.
In our comparison, we cross-evaluate each trained model’s performance on the test sets of both VWW and Wake Vision. Results on the Wake Vision test set can be found in Fig. 3, while the results for the VWW test set can be found in Fig. 5 alongside the complete data in Appendix A.
For the majority of models, the one trained on the Wake Vision (Quality) set achieves a higher accuracy on both the VWW and Wake Vision test set, which supports our claim that Wake Vision is a direct improvement over VWW. In some models, the VWW trained variant performs better on the VWW test set. This is only the case for the smallest models in our evaluation suite, which we believe may be due to these models being too small to generalize past the slight domain shift between the VWW and Wake Vision test sets. We additionally see that the accuracy discrepancy between identical models is often highest on Wake Vision’s Test set, further suggesting that the new test set is more challenging.
The boost in accuracy achieved by using the Wake Vision (Quality) Training set can be traded off for smaller model architectures (in some cases 3X fewer MACs) while preserving test accuracy. Accuracy and performance metrics, such as latency or model size, are in constant tension in TinyML models. Therefore, any boost in accuracy that does not impact the model’s architecture enables further efficiency.
4.4 Impact of Dataset Quality & Size in TinyML
Understanding the relationship between dataset quality and size is important for training efficient models in real-world scenarios. While gathering large amounts of data has become easier, ensuring high-label quality remains costly and time-consuming. This creates an important trade-off: is it better to have a smaller dataset with high-quality labels or a larger dataset with more label noise? The answer likely depends on the capacity of the model and the specific task.
@@ 4.4 Impact of Dataset Quality & Size in TinyML
We investigate the impact of dataset quality and size on Resnet [14] style models of varying capacity. We measure quality by the approximate rate of label errors. The Wake Vision (Quality) training set has an estimated error rate of around 7%. We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability. Appendix E gives the derivation for the flip probability. We also sweep the dataset size by taking a slice of Wake Vision (Quality). Each model is trained for 50,000 steps on 224x224x3 images using AdamW [26], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
Figure 4 shows that smaller models (leftmost figure) are more sensitive to higher error rates than large models, similar to prior findings that model capacity can help memorize and ignore errors and outliers [40, 9, 13]. The lowest accuracy drop in the smallest model is 1. 3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model. In contrast, large models benefit more from big datasets and start to overfit on the smaller slices of the training data [18]. We also see the large slice of the 15% error rate dataset outperforming a smaller slice of the 7% error rate dataset.
@@
Figure 4: Impact of dataset error rate and size on models of varying capacity. The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training. The error % indicates the expected rate of incorrectly labeled samples.
4.5 Fine-Grained Benchmark Evaluation
Although aggregate metrics like overall accuracy or F1 score provide a high-level view of model performance, they mask critical failure modes and biases that emerge in real-world scenarios. For this reason, we included fine-grained benchmarks in Wake Vision (Section 3.3). Fine-grained evaluation across demographics, environmental conditions, and visual contexts helps us develop reliable TinyML Computer Vision systems for bespoke use cases. This is particularly important for person detection models deployed in real-world applications, where failures on certain subgroups or conditions could have serious implications for fairness and safety.
In Table 5, we compare the models from Sec. 4.2, trained on Wake Vision and VWW, across various scenarios to understand their relative strengths and limitations. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66. Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82.
@@ 4.5 Fine-Grained Benchmark Evaluation
Gender Female $684$ $2157$ $0.93$ $0.89$
Male $1310$ $3918$ $0.91$ $0.88$
Unknown $1612$ $4940$ $0.77$ $0.78$
@@
Unknown $1299$ $3837$ $0.71$ $0.74$
Distance Near $5457$ 16 3331633316\,33316 333 $0.91$ $0.89$
Medium $2213$ $6876$ $0.85$ $0.84$
Far $398$ $1140$ $0.59$ $0.67$
Lighting Dark $3255$ $9420$ $0.85$ $0.81$
Normal 14 3151431514\,31514 315 43 0104301043\,01043 010 $0.82$ $0.82$
Bright $1012$ $3332$ $0.82$ $0.82$
@@
No Depiction $8583$ 25 8022580225\,80225 802 $0.87$ $0.85$
The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively. This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
Notably, a VWW model only significantly outperforms a Wake Vision model on the far distance dataset at a score of 0.67 and 0.59, respectively. This is likely caused by a difference in how the two datasets are created. Where VWW filters out persons smaller than 0.5% of the image, this filter is set to 5% for Wake Vision. Therefore, the VWW dataset contains images with significantly smaller persons than the Wake Vision dataset, which can improve its performance on far away persons. We did experiment with setting the limit to 0.5% for Wake Vision however, we experienced that it disproportionately decreased the performance of models trained on the dataset in other settings.
@@ 5 Ethical Considerations
The Wake Vision dataset and benchmark suite are designed to advance TinyML research by providing data resources and enabling detailed evaluation of person detection systems. While this work has significant potential benefits, we acknowledge several important ethical considerations:
Intended Benefits
Wake Vision can improve the quality of person detection systems that serve beneficial societal purposes, for example, an occupancy detection system that conserves energy by turning off the lights/HVAC/TV when a person leaves. The focus on TinyML and on-device inference helps protect user privacy by avoiding cloud-based processing of sensitive data.
Potential Misuse
We recognize that person detection technology can be misappropriated for harmful purposes, including weapons targeting or mass surveillance systems. While we cannot completely prevent misuse, we implement several safeguards, including releasing only evaluation sets, not training sets, for our fine-grained benchmark datasets.
Data Rights and Privacy
@@
The images in Wake Vision are sourced from Flickr through Open Images [21, 19] under the CC-BY 2.0 license. While we have made extensive efforts to ensure proper licensing, we acknowledge the possibility that some images may have been uploaded without proper distribution rights. Furthermore, the dataset may contain potentially identifying biometric information, and given sufficient resources, individuals could potentially be identified. Due to the scale of the dataset, exhaustive manual verification for offensive content or privacy concerns is infeasible.
Fairness and Representation
Our benchmark suite is specifically designed to evaluate model fairness and robustness across different demographics and conditions. However, we acknowledge that these same demographic benchmarks could be misused to develop systems that classify gender and age without consent, potentially enabling privacy violations or discriminatory practices. We have carefully balanced these concerns against the importance of ensuring that person detection systems work equitably for all populations.
6 Conclusion
The lack of large, high-quality datasets has been a significant bottleneck in TinyML research. Through Wake Vision, we address this challenge by providing a comprehensive dataset and benchmark suite that advances the field in several key dimensions. Wake Vision surpasses the current state-of-the-art VWW dataset by offering  100x more images while maintaining ease of use and improving label accuracy. Our experiments show that Wake Vision serves as an effective drop-in replacement for VWW, consistently improving performance across various model architectures and training scenarios. Moreover, our investigation into dataset quality versus quantity trade-offs reveals important insights specific to TinyML systems. Unlike large models that can effectively learn from noisy data, we find that TinyML models are particularly sensitive to label quality because of their limited capacity. However, we also demonstrate that carefully designed training strategies—such as our approach of pre-training on Wake Vision (Large) followed by fine-tuning on Wake Vision (Quality)—can effectively leverage both quantity and quality to achieve superior performance. Finally, the Wake Vision benchmark suite, with its fine-grained evaluation sets, establishes new standards to evaluate the robustness of the TinyML model in demographics, environmental conditions, and visual contexts. These benchmarks enable researchers to identify and address potential biases and failure modes that could otherwise go undetected using traditional evaluation metrics.
Appendix A Additional Benchmark Results
Table Tab. 6 provides input shape, Flash, RAM, and MACs for each model presented in section 4.2 as well as the mean and standard deviation of test accuracies for both Wake Vision and VWW datasets. RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”. Links pointing to the original models used for training are also present for each model family involved in the experiment.
Figure Fig. 5 shows cross-evaluation results on the Visual Wake Words test set. The majority of models trained on the Wake Vision dataset obtain better test accuracy on the VWW than the ones trained on the VWW. Only the smallest models do not achieve a higher test accuracy when trained on the Wake Vision, likely due to the lack of enough resources to overcome the slight domain shift between the two datasets.
Model input RAM Flash MACs Test Set Training Set
.tflite volume [kiB] [kiB] [MM] WV (Qual.) VWW
MobileNetV2  [^1] https://keras.io/api/applications/mobilenet/[33]
MobileNetV2_0.25 (224,224,3) 1,244.5 410.55 36,453,732 WV $84.9\pm 0.11$ $83.8\pm 0.23$
VWW $88.6\pm 0.17$ $88.3\pm 0.29$
MCUNet  [^2] https://github.com/mit-han-lab/mcunet/blob/master/mcunet/model_zoo.py[24]
10fps_vww (64,64,3) 168.5 533.84 5,998,334 WV $81.7\pm 0.28$ $77.6\pm 0.31$
VWW $81\pm 0.12$ $80.8\pm 0.1$
5fps_vww (80,80,3) 226.5 624.55 11,645,502 WV $82.9\pm 0.29$ $79.6\pm 0.26$
VWW $83.2\pm 0.57$ $82.6\pm 0.16$
320kb-1mb_vww (144,144,3) 393 923.76 56,022,934 WV $85.6\pm 0.34$ $82.4\pm 0.62$
VWW $86.5\pm 1.01$ $86\pm 0.25$
Micronets [^3] https://github.com/ARM-software/ML-zoo/tree/master/models/visual_wake_words[2]
vww2_50_50_INT8 (50,50,1) 71.50 225.54 3,167,382 WV $71.9\pm 0.67$ $66.2\pm 0.88$
VWW $70.8\pm 0.68$ $70.6\pm 0.64$
vww3_128_128_INT8 (128,128,1) 137.50 463.73 22,690,291 WV $77.8\pm 0.56$ $72.6\pm 1.01$
VWW $78.3\pm 0.91$ $77.6\pm 1.13$
vww4_128_128_INT8 (128,128,1) 123.50 417.03 18,963,302 WV $77.9\pm 0.6$ $71.3\pm 1.03$
VWW $78.4\pm 0.51$ $76.5\pm 0.37$
ColabNAS444 https://github.com/AndreaMattiaGaravagno/ColabNAS/tree/main/models [11]
k_2_c_3 (50,50,3) 18.5 7.66 250,256 WV $70.6\pm 0.96$ $69.3\pm 0.97$
VWW $65.6\pm 0.66$ $70.7\pm 0.08$
k_4_c_5 (50,50,3) 22 18.49 688,790 WV $75.7\pm 0.18$ $74\pm 0.23$
VWW $69.9\pm 0.26$ $75.5\pm 0,64$
k_8_c_5 (50,50,3) 32.5 44.56 2,135,476 WV $77.3\pm 0.37$ $75\pm 0.15$
VWW $73\pm 0.91$ $77.3\pm 0.57$
Table 6: Raw data of reconstructed models.
Figure 5: Cross-evaluation results on the Visual Wake Words test set
Appendix B Code Repository